### v2

# Challenging VLMs’ Structural Spatial Intelligence through Complex Reasoning Tasks

Abstract
Large Language Models (LLMs) have undergone rapid progress, largely attributed to reinforcement learning on complex reasoning tasks. In contrast, while spatial intelligence is fundamental for Vision-Language Models (VLMs) in real-world interaction, the systematic study of their complex spatial reasoning remains underexplored. To bridge this gap, we introduce SIRI-Bench, a benchmark designed to evaluate VLMs’ structural spatial intelligence through spatial-grounded reasoning tasks. SIRI-Bench comprises 9,000 video-question-answer triplets, where each problem is embedded in a realistic 3D scene. The benchmark is carefully designed so that solving each problem requires both spatial comprehension and structural reasoning. To facilitate large-scale data synthesis, we develop an Automatic Scene Creation Engine that employs collaborative LLM agents to translate abstract mathematical problems into faithful 3D scenes. Experimental results reveal that state-of-the-art VLMs struggle significantly on SIRI-Bench, underscoring the challenge of structural spatial reasoning. We hope that our study will bring researchers’ attention to spatially grounded reasoning and advance VLMs in visual problem-solving.
1 Introduction
Recent advancements in Large Language Models (LLMs) and Vision-Language Models (VLMs) have driven the emergence of structural reasoning, enabling strong performance on complex tasks (Wei et al., 2022; Lu et al., 2025; Fan et al., 2025; Jaech et al., 2024; Guo et al., 2025). Through extensive training on mathematics and programming, these models exhibit remarkable generalization across diverse, intricate challenges (Lewkowycz et al., 2022; Achiam et al., 2023; Chen et al., 2024b; Saab et al., 2024; Chen et al., 2024c; 2021b; Li et al., 2022).
With growing interest in complex problem solving, numerous benchmarks have been proposed to assess the reasoning ability of LLMs and VLMs (Hendrycks et al., 2021; Liu et al., 2024; Chen et al., 2024d; Fu et al., 2024). However, most of them focus on abstract tasks such as algebra, program synthesis, or geometric reasoning (Lu et al., 2023; Wang et al., 2024a; Chen et al., 2021a; Zhang et al., 2024a; Amini et al., 2019; Zhang et al., 2024b), largely overlooking visual-based reasoning tasks that are crucial for real-world interaction, particularly, the spatial intelligence.
Spatial intelligence refers to the capacity to interpret spatial information (Gardner, 2011), encompassing skills such as perceiving shapes, understanding transformations, and applying spatial knowledge (Yang et al., 2024c). This capability is essential for the VLMs and downstream real-world deployment (Brohan et al., 2023; O’Neill et al., 2024; Mangalam et al., 2023; Chandrasegaran et al., 2024). While notable studies have explored spatial intelligence (Yang et al., 2024c; Du et al., 2024; Man et al., 2024; Cai et al., 2025), they have largely focused on tasks like distance estimation or layout understanding, which tend to emphasize intuitive intelligence. By contrast, relatively little attention has been devoted to the complex reasoning levels that are essential for solving spatially grounded problems, a capacity we refer to as structured spatial intelligence.
To address this limitation, we introduce the Spatial Intelligence ReasonIng Benchmark (SIRI-Bench), a dataset of 9,078 video–question–answer triplets specifically designed to evaluate VLMs’ structural spatial intelligence through complex reasoning tasks. Following the previous paradigm of mathematical structural reasoning (Zhang et al., 2024b; Lu et al., 2023; Wang et al., 2024a; Yue et al., 2024; the University of Utah, 2024; Lightman et al., 2023), we construct SIRI-Bench based on solid geometry problems. Unlike conventional text or 2D diagram settings, where conditions are explicitly described (fig. 1 upper-left), SIRI-Bench encodes them in 3D scenes and rendered as videos (fig. 1 lower-left). In this representation, key conditions and spatial relations are implicitly embedded, requiring models to interpret and infer them from spatial cues. This design ensures that both spatial perception and structural reasoning are indispensable for solving the questions, providing a systematic and challenging benchmark for assessing VLMs on spatially grounded reasoning.
Figure 1: Spatial Intelligence with Complex Reasoning. Conventional benchmarks focus on text-grounded reasoning (upper-left), where reasoning is limited to text alone. In contrast, this work introduces spatial-grounded reasoning (lower-left), where key conditions are implicitly embedded in realistic 3D scenes and presented as videos. Solving these problems requires interleaved textual reasoning and spatial perception (right col.), challenging VLMs’ spatial reasoning ability.
Constructing such a dataset at scale entails significant challenges, as manually annotating is highly labor-intensive and demands substantial expertise. To overcome this, we develop an Automatic Scene Creation Engine that translates 3D geometry problems into realistic 3D scenes and renders them into videos. The engine leverages multiple specialized LLM agents within a sophisticated workflow, enabling the automatic computation of object dimensions, Blender scripting, and video rendering. It also adapts problem text to ensure that solving requires genuine spatial reasoning rather than textual shortcuts. As demonstrated by the examples in fig. 3, this engine can process diverse geometry problems and produce faithful 3D scenes, supporting large-scale benchmark creation.
We conduct extensive experiments on SIRI-Bench with a range of popular VLMs. Results show that even SOTA models fail on over 50% of the problems, with prediction errors exceeding 80% compared to ground truth. Remarkably, when key mathematical conditions (e.g. object dimensions and geometric types) are explicitly provided in text, performance improves by more than twofold, indicating models’ inability to extract them from video. Further comparisons with human participants and qualitative analyses confirm that current VLMs fall short on SIRI-Bench, underscoring their limitations in spatially grounded reasoning and the value of our benchmark.
The major contributions of this paper are summarized as follows: (1) We introduce the SIRI-Bench, a benchmark designed to investigate VLMs’ structural spatial intelligence through complex reasoning tasks. By representing problems through video-based 3D representations, SIRI-Bench establishes a novel framework for evaluating VLMs’ structural reasoning capability grounded in spatial comprehension. (2) We develop an Automatic Scene Creation Engine that translates 3D geometry problems into realistic 3D scenes. By leveraging specialized LLM agents with a structured workflow, the engine can produce faithful 3D scenes and significantly reduce the cost of large-scale data generation. (3) We benchmark the performance of state-of-the-art VLMs on SIRI-Bench, and find that they struggle to extract critical spatial information from visual inputs when solving complex reasoning tasks, revealing key limitations in spatially grounded reasoning.
2 Related Work
2.1 Spatial intelligence
Spatial intelligence, a core component of cognition, was introduced by Davison (2018) as an extension of visual SLAM (Simultaneous Localization and Mapping). Following this, Chen et al. (2024a) proposed SpatialVLM, trained on 3D visual QA, achieving improved spatial estimation, while Yang et al. (2024c) defined visual-spatial intelligence as the ability to process visual information in 3D space. Other studies have also introduced benchmarks targeting spatial intelligenceWu et al. (2023b); Yang et al. (2024c); Du et al. (2024); Cai et al. (2025). Despite these advances, current VLMs largely remain at the level of intuitive intelligence, such as perception, whereas this study focuses on structured spatial intelligence requiring complex reasoning.
2.2 Complex Reasoning
Complex reasoning involves multi-step inference, abstract thinking, and knowledge integration. Prior work shows that LLMs are capable of such reasoning, mainly through reinforcement learning on mathematics. (Shao et al., 2024; Lu et al., 2025; Guo et al., 2025; Yang et al., 2025). Their applications span mathematics (Lewkowycz et al., 2022; Hendrycks et al., 2021; Achiam et al., 2023; Jaech et al., 2024), medical diagnosis (Chen et al., 2024b; Saab et al., 2024; Chen et al., 2024c), and programing  (Chen et al., 2021b; Li et al., 2022; Jaech et al., 2024; Guo et al., 2025), etc. However, LLMs show limited capability in visual-language reasoning (Małkiński et al., 2024; Chia et al., 2024; Ghosal et al., 2024). Although VLMs show proficiency in conventional video question-answering tasks (Wang et al., 2024b; Team, 2024a; Lin et al., 2023; Maaz et al., 2023), their capacity for integrating complex reasoning with spatial intelligence remains underdeveloped. To address this limitation, this paper represents math problems through video-based 3D scenes that jointly demand both spatial understanding and structural reasoning.
2.3 VLM Reasoning Benchmark
A number of benchmarks have been proposed to assess VLMs’ visual reasoning, often by reducing linguistic biases and enforcing reliance on visual input (Johnson et al., 2017; Goyal et al., 2017; Hudson & Manning, 2019; Liu et al., 2024; Chen et al., 2024d; Fu et al., 2024). Others, such as MathVista (Lu et al., 2023), MATH-Vision (Wang et al., 2024a) , GeoQA (Chen et al., 2021a), Geoeval (Zhang et al., 2024a), Mathqa (Amini et al., 2019) and MathVerse (Zhang et al., 2024b) , target symbolic and geometric reasoning through plane geometry, but remain abstract and detached from real-world contexts. In contrast, our benchmark is grounded in realistic 3D environments with precise geometric properties (e.g., angles, distances) and requires multi-step inference that integrates spatial perception with procedural reasoning.
3 SIRI-Bench
Figure 2: The Transformation Process from an original math problem to a 3D Spatial Representation. The given math problem is decomposed into five components and processed individually. First, the main entity’s dimensions are solved, and corresponding bpy code generates the 3D scene, later rendered as a video. Second, problem conditions are refined by removing information meant to be inferred from the scene, and node indices are replaced with color markers. Finally, the answer is adjusted for scaling effects to produce the final answer.
To evaluate VLMs’ spatial intelligence in complex reasoning, we introduce SIRI-Bench, a benchmark built from solid geometry problems and cast as video-based question–answering problems. For each problem, we elaborately design the 3D spatial representation***The term 3D spatial representation in this paper refers to representing math problems as 3D scenes. It should not be confused with learnable 3D parameters, such as those used in 3D Gaussian splatting. that embeds key mathematical conditions into realistic 3D scenes, requiring VLMs to extract relevant information through spatial perception and reasoning.
3.1 Data Collection
In line with prior work, our benchmark is built upon math problems, as they inherently involve multi-step and structural reasoning (Zhang et al., 2024b; Lu et al., 2023; Wang et al., 2024a; Yue et al., 2024; the University of Utah, 2024; Lightman et al., 2023). The key distinction is that we focus on 3D geometry problems rather than algebra or plane geometry problems, and transform them into realistic 3D scenes for spatial reasoning.
Specifically, SIRI-Bench is constructed from publicly available 3D geometry math problems collected from online educational resources. All problems are translated into English and span difficulty levels from middle to high school. To ensure consistency, we remove low-quality problems, proof-based questions, and other non-numerical items, leaving only problems that require explicit value computation. These are then converted into open-ended problem-solving.
3.2 3D Spatial Representation
This section introduces the 3D spatial representation, which encodes geometry problems as realistic 3D scenes rendered into videos. Compared to traditional abstract representations such as textual descriptions or 2D diagrams, our representation offers a closer approximation of real-world scenarios and requires VLMs to extract information from spatial cues rather than textual descriptions.
Specifically, an original geometry problem can be decomposed into five components: (1) Illustration Diagram, (2) Geometry’s Category, (3) Primary Conditions, (4) Auxiliary Conditions, and (5) Final Question, each of which is transformed into the 3D spatial representation through distinct processing steps, thereby depicting the problem anew within the 3D scene.
For the Illustration Diagram, it is typically used to visually represent the problem. In our approach, we discard the diagram and displaying the problem through a 3D realistic video.
For the Geometry’s Category, the main geometric entity is directly presented in the 3D scene. Instead of naming it explicitly (e.g., “the pyramid” or “the cube”), the text only refers to it as “the object” leaving the model to recognize its shape from the visual cue. To avoid information leakage from vertex indices (e.g., “PABCD” indicating a quadrangular pyramid), we replace letter-based indices with color-based indices, with colored points placed in the scene to ensure unambiguous references. For example, the vertex ‘$S$’ in fig. 2 is replaced with the $P_{orange}$’.
The primary geometric conditions refer to the attributes and constraints of the main geometric entity, such as its dimensions and spatial relations. Instead of appearing in the text, these conditions are converted into concrete geometric sizes or spatial positions and directly visualized in the 3D scene, enabling them to be inferred through observation.
The auxiliary conditions, which provide necessary but non-central information, are preserved in textual form due to their diverse and complex nature. Similarly, the final question, specifying what ultimately needs to be solved, is also kept in textual format.
Our approach is conceptually aligned with MathVerse (Zhang et al., 2024b), which embeds math conditions into 2D diagrams. However, our method goes beyond abstract diagrams by encoding key information in 3D scenes, bringing the problems closer to real-world scenarios. This requires not just optical character recognition (OCR) but genuine spatial perception and problem-solving.
3.3 Automatic Scene Creation Engine
Manually converting each problem into a 3D scene is extremely labor-intensive. To facilitate large-scale 3D scene generation, this section introduces an Automatic Scene Creation Engine based on a multi-agent system (Wu et al., 2023a; Hong et al., 2023; Wei et al., 2024; Yang et al., 2024d; Yu et al., 2025). The workflow is shown in fig. 2.
First, the engine begins by parsing the original 3D geometry problem and identifying the main geometric entity. A math-specialist LLM agent (Qwen-Math (Yang et al., 2024b), known for its strong capability on mathematics) is employed to solve for all key conditions necessary to fully define the main geometric entity. After that, all dimensions are scaled to fit appropriately within a unified camera configuration, with sizes ranging from 0.5 to 2m. These parameters are passed to a code-specialist agent (Qwen-LM (Yang et al., 2024a)), which generates bpy scripts(Blender Online Community, 2022) to construct the scene and insert colored vertices (see section 3.2). Second, we refine the problem descriptions to retain only essential information while minimizing textual mathematics. This is achieved by an LLM agent (Qwen-LM (Yang et al., 2024a)), which removes key geometric conditions of the main entity. As discussed in section 3.2, this forces VLMs to rely on spatial comprehension rather than textual cues. Third, the answers are converted from symbolic expressions to numerical values by an LLM agent, allowing for direct measurement of prediction errors and analysis of their distribution. The agent also adjusts values to account for geometric scaling—linear for lengths, quadratic for areas, and cubic for volumes.
By elaborately designing the specialized agents and their workflow, our Automatic Scene Creation Engine is able to transform any 3D geometry problem into a faithful 3D scene. Using this pipeline, we constructed a dataset of 9078 samples for SIRI-Bench, after manually filtering out a few clear conversion errors. While moderate in size, the dataset demonstrates the engine’s scalability for generating larger datasets in the future.
Figure 3: Data Samples in SIRI-Bench. This figure presents several samples from our SIRI-Bench dataset, along with their original questions and intermediate steps. As can be consistently observed, our data generation engine accurately solves for geometric conditions, replaces vertex indices, processes textual conditions, and computes numerical answers, demonstrating its reliability.
3.4 Other Details
[…]
3.5 Data Samples in SIRI-Bench
In fig. 3, we present illustrative samples from SIRI-Bench, including original questions, final questions and intermediate steps. As shown, our data engine effectively conceals mathematical conditions that must be inferred from video while retaining auxiliary information, and accurately replaces vertex indices with spatially aligned references. Intermediate results further confirm the engine’s ability to compute object dimensions, even in complex cases such as sample #000010. Comparisons between original and final answers also validate precise numerical computation with scaling adjustments. Overall, these visualizations demonstrate the reliability of our Automatic Scene Creation Engine in generating faithful 3D scenes from diverse geometry problems.
Figure 4: Performance of Existing VLMs. This figure shows the error distributions across seven intervals ranging from 0% to 200% for all baseline methods on the SIRI-Bench. A higher concentration of errors in the lower intervals (i.e. brighter colors) indicates better performance in problem-solving. The method labeled ‘Textual Rep.’ refers to an LLM that accesses full mathematical conditions through textual descriptions rather than videos of 3D scenes. Overall, the results reveal the limitations of current VLMs in spatial grounded reasoning.
4 Experiment
Figure 5: Ablation on Problem Representation. This figure compares the accuracy of two sibling models using textual representation versus 3D spatial representation as input. Three columns depict three pairs of sibling LLMs/VLMs. This comparison disentangles structural reasoning from spatial perception, revealing that existing VLMs struggle to effectively extract spatial information when solving complex visual problems.
4.1 Experimental Setup
Evaluation Metrics. Our evaluation aims to jointly assess both spatial estimation accuracy and mathematical reasoning correctness. To achieve this, we compute the relative error between the predicted numerical answer and the ground-truth value, which can be expressed as: $\frac{{}|\hat{y}-y|}{y}\times 100\%,$ where $\hat{y}$ denotes the predicted numerical value and $y$ denotes the ground-truth numerical value. To better analyze performance, we categorize the relative error into seven intervals (0–20%, 20–40%, 40–80%, 80–100%, 100–200%, and >>200%) and examine the distribution of predictions. Strong models concentrate in lower-error intervals, whereas weaker models show higher errors.
Baselines. We evaluated a variety of VLMs, encompassing models of different sizes, both open/close-source models. They are***The specific versions we used is o3-2025-04-16, gpt-4o-2024-08-06, gpt-4-turbo-2024-04-09, qwen-vl-max-2025-08-13, qwen-max-2024-09-19, doubao-seed-1-6-vision-250815, doubao-1.5-vision-pro-250328: OpenAI: OpenAI-o3, GPT-4o, GPT-4v OpenAI (2023b; 2024; a). Qwen: Qwen2.5-VL-7B, 32B, 72B, QVQ-72B, and Qwen-VL-Max (Yang et al., 2024a; Team, 2024b; c). Doubao: Doubao-1.6-vision, Doubao-1.5-vision (Team, 2025b; a). InternVL: InternVL2.5-8B, 26B (Chen et al., 2024e). LLaVA-Next-Video: LLaVA-NV-7B, 34B (Zhang et al., 2024c). Others: SpaceQwen-3B (Chen et al., 2024a). LLM: Doubao-1.6 (Textual Rep.) (Team, 2025b), which accesses the full mathematical conditions through text rather than videos. We append it to disentangle structural reasoning from spatial perception.
4.2 Overall Performance
The overall performance of existing VLMs is summarized in fig. 4, which presents the distribution of prediction errors across seven intervals ranging from 0% to over 200%. A higher concentration of predictions in the lower error intervals indicates greater accuracy.
All evaluated VLMs perform poorly on SIRI-Bench. The best method, Doubao-1.6-Vision (Team, 2025b), achieves only 31% of predictions within a 0–40% error margin. This performance is consistent with its strong results on other multimodal reasoning benchmarks (Wang et al., 2024a; Lu et al., 2023). Nevertheless, over 33% of its predictions exceed 100% error, underscoring its substantial limitations. (Doubao-1.6-Textual-Rep. is discussed separately in section 4.3.) There is no evident correlation between performance and parameter size. For instance, Qwen2.5-VL improves from 3B to 7B, but InternVL2.5 declines when scaling from 7B to 26B. This implies that enhancing the spatial grounded reasoning ability may need other strategies beyond simply scaling up. Another noteworthy case is SpaceQwen-3B (Chen et al., 2024a), a small model fine-tuned for spatial perception. It achieves 37% of predictions under 60% error, showing competitive performance relative to larger models. However, its accuracy remains unsatisfactory, likely due to both its limited scale and a fine-tuning focus on intuitive perception rather than structural reasoning.
We hypothesize two main reasons for their poor performance: (1) models struggle to estimate object dimensions and spatial relationships since they are not explicitly provided; (2) models are prone to errors in mathematical reasoning and calculation. We will further analyze these issues in later sections. In summary, the consistent failure across models underscores the limitations of current VLMs in structural spatial reasoning.
Figure 6: Ablation on Deep-Thinking. This figure compares the performance with or without Deep-Thinking. The left two columns use 3D spatial representation and the right two use textual representation.
Figure 7: Error Distribution by Problem Type. The figure shows the success rate of five baseline models across five types of problems, where success is define relative error below 40%.
4.3 Ablation on Input Representation
As shown in fig. 4, the doubao-1.6-Textual-Rep. substantially outperforms all VLM baselines. This method is an LLM model that inputs textual representations and accesses mathematical conditions through text. To further investigate this phenomenon, we compare three pairs of sibling models: one LLM using text input (i.e. textual representation) and one VLM using video-text input (i.e. 3D spatial representation). The only difference lies in whether spatial information is directly presented in text or must be derived from video. This ablation disentangles structural reasoning from spatial perception, isolating the source of performance gaps.
The results in fig. 5 demonstrate a consistent trend: text-input models clearly outperform their video-input counterparts. For example, the proportion of predictions within 20% error nearly doubled when switching from video to text input (18.4%, 16.4%, 12.7% vs. 89.2%, 64.0%, 28.9%). These findings indicate that while SOTA LLMs can handle complex reasoning when conditions are provided in language, VLMs fail to do the same on the visual level, as they struggle to extract sufficient spatial information from video. Moreover, the strong performance of Doubao-1.6 validates the effectiveness of SIRI-Bench, as most questions are solvable with full textual conditions.
4.4 Ablation on Deep-Thinking
We conduct an ablation study to examine the effect of Deep-Thinking, which refers to an explicit deep reasoning phase before producing the final answer. As shown in fig. 7, when the input is 3D spatial representation, disabling Deep-Thinking for Doubao-1.6 reduces performance from from 31% to 25% (prediction error below 40%). In contrast, the impact is minimal with textual input, likely because direct answering is sufficient for most problems. These results suggest that deep reasoning plays a more critical role in spatial grounded reasoning.
4.5 Error Distribution
We analyze model performance across 5 problem types: angle, ratio, length, area, volume. We define a prediction as successful if its relative error is below 40% and report 5 corresponding success rates in fig. 7. There are two major observations. (1) All methods perform poorly on volume-related problems, with fewer than 20% of their predictions under the 40% error threshold. We suggest that errors in volume problems may stem from spatial estimation, as the deviations will be amplified cubically in the calculation. (2) Among the five baselines, Doubao-1.6-Vision achieves the best overall performance. Its advantage is mainly on angle and ratio problems, where the performance improves from 30% to 40% over the second-best model. Since solving angle and ratio questions primarily requires reasoning about spatial relations rather than absolute dimensions, this finding highlights Doubao-1.6’s strength in relational geometric reasoning and suggests room for improvement for other models in this aspect.
Figure 8: Analysis of Reasoning Mistakes and Their Impact. The lower part shows the proportion of VLM reasoning responses exhibiting each of the three mistakes, where True means no mistake and False otherwise. The upper part illustrates the median prediction error associated with each condition, revealing how specific reasoning mistakes affect final performance.
4.6 Full Response Analysis
We further analyze VLMs’ full reasoning response by examining three types of mistake, mathematical calculation errors, erroneous assumptions, and geometric type misjudgments. Specifically, we provide VLMs’ full reasoning response to another LLM, which identifies whether any of these mistake occur. For geometric type mistake, the correct geometric type is also provided. We then visualize and study how these mistakes correlates with the error distribution in fig. 8.
Across all three categories, the occurrence of any mistake consistently leads to higher prediction errors. The most notable case is erroneous assumptions, where the median error increases from 0.56 to 0.90 for Qwen-VL-Max. Moreover, VLMs often misidentify the geometric type. For instance, GPT-4o exhibits such mistakes in nearly 79% of the problems. These results indicate that failures in geometric type recognition and the tendency to introduce unsupported assumptions are key factors limiting VLMs’ reasoning accuracy.
4.7 Comparison with Human Performance
Figure 9: Comparison with Human Performance. This figure compares the performance of human participants (the 1st bar) with four VLMs on a subset of the SIRI-Bench, showing that current VLMs are still far from matching human performance.
In this part, we investigate how far VLMs are from matching human performance. We randomly select seven samples from SIRI-Bench and ask participants to solve them without additional information. The average error distribution of human participants and VLMs is shown in fig. 9. As can be seen, humans participants successfully solve over 30% problems with errors within 20%. In contrast, none of the VLMs achieve this level of accuracy. When allowing for a 60% error margin, humans succeed on about 70% of the problems, far surpassing VLMs. Yet, some problems remain challenging even for humans, occasionally yielding errors above 100%, likely due to accumulated deviations in distance estimation. Nevertheless, the overall results demonstrate that current VLMs consistently fall short of human capability.
Figure 10: Display of GPT4o’s Full Answer. This figure displays the full response from GPT-4o on an example, with spatial interpretation errors, spatial perception errors, and the final incorrect result highlighted in red, blue, and orange, respectively. The visualization illustrates that GPT-4o struggles to extract problem-solving conditions from visual inputs, and tends to make unjustified assumptions that ultimately lead to incorrect answers.
4.8 Case Study
Figure 10 presents a case study of GPT-4o’s reasoning process, with key errors highlighted. This case illustrates potential factors contributing to GPT-4o’s poor performance on SIRI-Bench. Specifically, the question asks for the area of the cross-section formed by the plane defined by points D, P, and Q intersecting a cube. GPT-4o oversimplifies the problem by assuming that three points define a triangular cross-section (DPQ), whereas the actual cross-section is a rectangle. Additionally, without effectively extracting spatial cues from the video, it further assumes the triangle is right-angled and estimates the cube’s size as unit-length, ultimately leading to an incorrect answer.
Overall, this example highlights GPT-4o’s difficulty in interpreting geometric constraints and its tendency to make unjustified assumptions when visual cues are implicit. Rather than leveraging the video to resolve ambiguity, GPT-4o often proceeds with incomplete reasoning, revealing a notable problem-solving limitation in the visual domain.
5 Conclusion
This paper presents SIRI-Bench, a benchmark designed to assess the structural spatial intelligence of VLMs through 3D mathematical reasoning problems. SIRI-Bench consists of 9078 video-question-answer triplets, where each question is formulated within a realistic 3D scene captured in video format. Solving these questions requires models to extract, interpret, and perform reasoning based on spatial cues, offering a novel and rigorous evaluation of VLMs’ spatial grounded reasoning ability. To construct this dataset efficiently, we develop an Automatic Scene Creation Engine that converts abstract math problems into realistic 3D scenes and renders corresponding videos. The engine coordinates multiple specialized LLM agents within a structured workflow, ensuring reliable scene generation while significantly reducing the cost of large-scale data production. Experiments on SIRI-Bench demonstrate that existing VLMs consistently fail to solve more than 50% problems when spatial information must be inferred from video. Their performance not only falls far behind human level but also lags significantly behind their text-only counterparts, which are provided with explicit mathematical conditions in text form. Further experiments reveal that existing VLMs over-rely on textual cues and struggle to actively and accurately extract key spatial information—such as object types, geometric dimensions, and spatial relations—from visual input. These findings highlight a fundamental limitation of current VLMs and demonstrate the value of SIRI-Bench***Upon acceptance, we will release the code and dataset for public access.
Ethics statement
Our work is conducted in accordance with ethical research practices and legal requirements. All original mathematical problems used in SIRI-Bench are licensed for research use, and the 3D scenes are generated based on the publicly available 3D-FRONT dataset, for which we have obtained proper authorization. The data generation process involves only synthetic scenes and automatically rendered videos, without collecting or inferring any personally identifiable information (PII). We emphasize that SIRI-Bench is designed solely for academic research on spatial reasoning and model evaluation. It should not be misused for surveillance, discriminatory applications, or other purposes that could harm individuals or communities.
Reproducibility Statement
We provide comprehensive details of our dataset construction pipeline. All prompts used in data generation and experiments are included in the Appendix to enable precise replication. We also describe our experimental settings, evaluation protocols, and ablation study configurations in detail. We will publicly release the SIRI-Bench dataset, along with the associated code and generation scripts, upon acceptance. This ensures that all results presented in the paper can be independently verified and extended by the research community.
LLM Clarification
All uses of Large Language Models (LLMs) in this work are explicitly documented in the paper. During dataset construction, they were used solely as auxiliary tools for repetitive tasks such as condition text editing and dimension solving, always under human supervision. Crucially, all core scientific contributions—including the formulation of the research problem, the design of the SIRI-Bench dataset, the development of the scene generation pipeline, and the experimental design and analysis—were conceived and executed by the authors. The use of LLMs did not involve conceptualization, methodology design, or result interpretation, and did not reach a level that would warrant considering them as contributors.